Session: Feb 2005 (session 1)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Medicina (mauro at Sat Feb 19 02:14:57 2005) Success :
Onsala (Michael Lindqvist at Sat Feb 19 09:30:17 2005) Success :
No known problems. Some snow and rain.
Noto (P. Cassaro at Sat Feb 19 15:37:57 2005) Success :
Effelsberg (A Kraus at Mon Feb 21 12:25:32 2005) Minor failures : WIND
Observation was stopped due to snowfall between UT 1710-1742,
UT1753-1824, and UT1951 until the end.
Metsahovi (Ari Mujunen at Mon Feb 21 13:12:51 2005) Success :
No known problems, pcal turned off manually, some light snow during the experiment, so weather was not terrific.
Jodrell_Bank (Alastair Gunn at Mon Feb 21 21:02:34 2005) Minor failures :
MkII was used.
Lost communication with Mk5 unit at about 07:58 UT (day 49). Required reboots; back on scan beginning 09:23 UT. Also seems to have stopped recording at 22:15 UT (scan 302) due to unknown Mk5 problem. No other known problems. Rain through much of experiment.
Cambridge (Alastair Gunn at Mon Feb 21 21:03:11 2005) Success :
No known problems.
Robledo (Cristina Garcia Miro at Tue Feb 22 15:33:06 2005) Minor failures : RECO
FS version 9.6.9.
At 17:49:40 low tape sensor stopped the tape unexpectedly. Tape was positioned manually for following source. Not enough time was provided to swap tapes (only 15 minutes, 13 minutes were needed to rewind the tape). Resume recording at 19:49UT.
Tsys data provided in observation log.
Not able to provide gps-fmout data due to a communication problem with the frequency counter.
> Weather condition: clear
> Station Time delay: -1,5451 us
> Recorder used: # 1
> Number of Tapes: 2
> Label(/s): hst00105,d43a - evnt0493, 7ff9
> Calibration at zenith:
> K-Band (LCP) System Temp: 43,26
> K-Band (LCP) Y-factor: 8,7
> K-Band (LCP) Noise Diode: 19,1
> K-Band (RCP) System Temp: 44,5
> K-Band (RCP) Y-factor: 8,4
> K-Band (RCP) Noise Diode: 15,9